OpenStudy (anonymous):

The vertex form of the equation of a parabola is y = (x - 5)2 + 16. What is the standard form of the equation? A. y = x2 - 10x + 41 B. y = x2 + 10x + 16 C. y = x2 + x + 8 D. y = 3x2 - 10x + 41

OpenStudy (anonymous):

\[ y = (x - 5)^2 + 16\\ y=(x-5)(x-5)+16\] etc

OpenStudy (anonymous):

multiply out, combine like terms

OpenStudy (anonymous):

A. y = x2 - 10x + 41 B. y = x2 + 10x + 16 C. y = x2 + x + 8 D. y = 3x2 - 10x + 41

OpenStudy (zzr0ck3r):

@DonJesusMedina do what @satellite73 said to get the answer

OpenStudy (anonymous):

is it D ? Please i need to be sure ?